6. Anthon received a commission of Php 214,000 for selling a condominium umit. If the rate of commission is 2.5%, how much is the price of the unit?
7. Archie bought a laptop and was charged Php 5,400 for a 12% VAT, how much is the apied tax rate?
8. Engr. Miguel was chargedPhp 18,560 as withholding a tax for his fees as a consultant to a certain project. If Engr. Miguel's consultant fee is Php 185,600, how much is the applied tax rate?
9. In 2014, Angelique flower shop sold a total of 15,000 dozen roses while 18,000 dozen roses were sold in the year 2015. What is the percent increase in the sale of the roses?
10. In a sirvey of 500 people, 32% said that they prefer to purchase Brand A ice cream than other brands. How many people prefer Brand A? How many peor prefer the other brands?
1
Expert's answer
2021-10-04T18:35:09-0400

6) Price of the unit"=100/2.5\times 214000=Php8560000"

7) applied tax"=12/100\times 5400=Php648"

8) applied tax rate"=18560/185600\times 100\%=10\%"

9) Percentage increase in the sale of roses "=\frac{18000-15000}{15000}\times 100\%=3000/15000\times 100\%=20\%"

10) People that prefer brand A "=32/100\times 500=160"

People that prefer other brands "=500-160=340"